OpenStudy (lacris):

if log10(x) + log10 (y) = 3 then xy = ?? please help with this solution, I will give you a shiny medal! and do show detailed instructions please :) :D Thank you

OpenStudy (luffingsails):

The important thing is that you have the same log base... So log10(x) + log10(y) can be rewritten as log10(xy)

OpenStudy (melbax14):

assuming the logs are base 10, i wont write the 10 for sake of clarity: logx + logy = 3 log(xy) = 3 raise both sides to power 10 (as this is the opposite of "taking logs", we get: xy = 10^3 so xy = 1000

OpenStudy (lacris):

So its just log(xy) =3 ? so does it connect to the rule with logs that addition is = or turned into multiplication? :D

OpenStudy (lacris):



OpenStudy (lacris):

as melbax said, it would be 1000 because 10^3 = 1000 right?

OpenStudy (lacris):

thank you both so much for helping me! I wish I could give you both medals!